# Revisiting In-context Learning Inference Circuit in Large Language Models

### v1

@@ Abstract
In-context Learning (ICL) is an emerging few-shot learning paradigm on Language Models (LMs) with inner mechanisms un-explored. There are already existing works describing the inner processing of ICL, while they struggle to capture all the inference phenomena in large language models. Therefore, this paper proposes a comprehensive circuit to model the inference dynamics and try to explain the observed phenomena of ICL. In detail, we divide ICL inference into 3 major operations: (1) Summarize: LMs encode every input text (demonstrations and queries) into linear representation in the hidden states with sufficient information to solve ICL tasks. (2) Semantics Merge: LMs merge the encoded representations of demonstrations with their corresponding label tokens to produce joint representations of labels and demonstrations. (3) Feature Retrieval and Copy: LMs search the joint representations similar to the query representation on a task subspace, and copy the searched representations into the query. Then, language model heads capture these copied label representations to a certain extent and decode them into predicted labels. The proposed inference circuit successfully captured many phenomena observed during the ICL process, making it a comprehensive and practical explanation of the ICL inference process. Moreover, ablation analysis by disabling the proposed steps seriously damages the ICL performance, suggesting the proposed inference circuit is a dominating mechanism. Additionally, we confirm and list some bypass mechanisms that solve ICL tasks in parallel with the proposed circuit.
@@ 1 Introduction
In-Context Learning (ICL) (Radford et al., 2019; Dong et al., 2022) is an emerging few-shot learning paradigm: given the demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ consisting of [input text]-[label token] pairs and a query $x_{q}$, Language Models (LMs) take the sequence $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$ [^1] In this paper, we denote tokenization as []delimited-[][$\cdot$][ ⋅ ]. as input and then predicts the label for $x_{q}$ by causal language modeling operation. Typically, the label tokens $y_{i}$ are preceded by and also predicted by forerunner tokens $s_{i}$ (e.g., the colon in “Label: ”). ICL has aroused widespread interest, but its underlying mechanism is still unclear.
@@
As shown in Fig. 1, we decompose ICL dynamics into 3 atomic operations on Transformer layers. Step 1: Summarize: LMs encode each input text $x_{i}$ into linear representations in the hidden state of its corresponding forerunner token $s_{i}$. Step 2: Semantics Merge: For demonstrations, LMs merge the encoded representations of $s_{i}$ with the hidden state of its corresponding label tokens $y_{i}$. Step 3: Feature Retrieval and Copy: LMs retrieve merged label representations $y_{1:k}$ from Step 2 similar to the query representation $s_{q}$ in a task-relevant subspace and then merge them with the query representation. Finally, LM heads predict the label for $x_{q}$ using the label-attached query representation $s_{q}$. Steps 2 and 3 form a typical induction circuit, which is a key mechanism of ICL but only examined in synthetic scenarios (Elhage et al., 2021; Singh et al., 2024b; Reddy, 2024).
@@
We empirically find evidence for the existence of each proposed step in LLMs, and conduct more fine-grained measurements to gain insights into some phenomena observed in ICL scenarios, such as (1) position bias: the prediction is more influenced by the latter demonstration (Zhao et al., 2021), (2) noise robustness: the prediction is not easy to be affected by demonstrations with noisy labels (Min et al., 2022), while larger models are less robust to noise (Wei et al., 2023), and (3) demonstration saturation: the improvements plateau when sufficient demonstrations are given (Agarwal et al., 2024; Bertsch et al., 2024), etc. (discussed in §5.3). Moreover, we find multiple bypass mechanisms for ICL with the help of residual connections, while the 3-phase dynamics remains dominant.
@@ 2.2 Experiment Settings
Models. We mainly conduct experiments on 4 modern LMs: Llama 3 (8B, 70B) (AI@Meta, 2024), and Falcon (7B, 40B) (Almazrouei et al., 2023). Unless specified, we report the results on Llama 3 70B, since its deep and narrow structure (80 layers, 64 heads) makes it easier to show hierarchical inference dynamics (discussed in §5.2). The results of other models can be found in Appendix F.2.
Datasets. We build ICL-formed test inputs from 6 sentence classification datasets: SST-2 (Socher et al., 2013), MR (Pang & Lee, 2005), Financial Phrasebank (Malo et al., 2014), SST-5 (Socher et al., 2013), TREC (Li & Roth, 2002; Hovy et al., 2001), AGNews (Zhang et al., 2015)specified, we report the average results on these datasets.
Others. Unless specified, we use $k=4$ demonstrations in ICL inputs. For each dataset, we randomly sample $n=512$ test data and assign one order-fixed demonstration sequence for each test sample. About the prompt templates, etc., please refer to Appendix A.1.
3 Step 1, Summarize: Linear Representations in Hidden States
@@
3.1 LLMs Summarize Input Text on Forerunner Tokens in Hidden States
We first study the existence of input text encoding in hidden states and then explain their linear separability and task relevance in §3.2. For each text-label pair $(x_{t},y_{t})$ (encoding target) sampled from the datasets, we prepend them with $k$ demonstrations, resulting in ICL-style inputs $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{t}][s_{t}][y_{t}]$. These inputs are then fed into an LM to extract the hidden states of a specific token in $[x_{t}][s_{t}][y_{t}]$ from each layer, serving as the ICL inner representations. To assess the quality of these representations as sentence representations, we use the sentence embedding of $x_{t}$ encoded by BGE M3 (Chen et al., 2024), a SotA encoder-only Transformer, as a “reference” representation and then calculate the mutual nearest-neighbor kernel alignment [^3] Intuitively, kernel alignment measures similarity between two representations toward the same datasets, and according to Huh et al. (2024), a higher cross-model kernel alignment usually means a better representation. (Huh et al., 2024) between these representations. See Appendix A.1 and A.2 for details.
@@ 3.1 LLMs Summarize Input Text on Forerunner Tokens in Hidden States
Perplexed Texts are Encoded Worse. We investigate the correlation between kernel alignment and the perplexity of encoding targets with different $k$. Fig. 2 (Right, upper) shows a negative correlation for $k=0$, that is, LMs generate poorer encodings for more complex input text when no demonstrations are given, which can be identified as an In-weight Learning (IWL) property of the inner text encoding. While, when demonstrations are given in context (Fig. 2 (Right, lower)), the negative correlation disappears, which suggests that LMs effectively encode more complex samples with the help of demonstrations in context. More discussion about the relationship of classification performance and perplexity is in Appendix E.
@@
3.2 Text Summarization is Linear and Task-Relevant but Position-biased
Figure 2: Input text encoding magnitudes (metricized by kernel alignment against feature encoded by encoder-structured models) of hidden states in various layers in ICL scenario (The controlled experiments are results between current 6 datasets and TEE (Mohammad et al., 2018)). Left: Encoding magnitudes on hidden states from various types of token. Middle [^5] Experiments of Fig. 2 (Middle) on Llama 3 70B do not involve results on AGNews.: Encoding magnitudes with different $k$ on the forerunner tokens. Right: Encoding magnitudes in layer 24 of Llama 3 70B against the causal language modeling loss of the input text with (upper) $k=0$ and (lower) $k=8$.
@@ 3.2 Text Summarization is Linear and Task-Relevant but Position-biased
4 Induction Circuits in Natural Language Models
Figure 4: The similarities of ICL hidden states in different positions on layer 24 between the same queries (Left) or two different queries (Right) (on SST-2).
This section mainly shows how LMs utilize the summarized linear text representations in induction circuits with a typical 2-step form (Singh et al., 2024b): Forerunner Token Heads merge the demonstration text representations in the forerunner token into their corresponding label tokens with a selectivity regarding the compatibility of demonstrations and label semantics. Induction Heads copy the information in the label representations similar to the query representations back to the query. Such operations are done on task-specific subspaces, enabling LMs to solve multiple tasks by multiplexing hidden spaces.
@@ 4.1 Step 2, Forerunner Token Head: Copy from Text Feature to Label Token
Text Representations are Copied without Selectivity. For each attention head, we extract the attention score $\alpha_{y_{t}\rightarrow s_{t}}$ from a label token $y_{t}$ (as attention query) to the corresponding forerunner token $s_{t}$ (as attention keys). We then mark the head with $\alpha_{y_{t}\rightarrow s_{t}}\geqslant 5/n_{t}$ ($n_{t}$: the length of tokens before $y_{t}$) as a Forerunner Token Head and count them in each layer. The results are shown in Fig. 5 (Middle, “Correct Label”), where the peak matches the copy period in Fig. 5 (Left). Moreover, to investigate the influence of the correctness of label tokens, we replace $y_{t}$ with a wrong label token [^7] For example, for a label space of “positive” and “negative”, if $y_{t}$ is “positive”, then replace it to “negative”., where the results in Fig. 5 (Middle, “Wrong Label”) are almost identical to the correct-label setting, suggesting that the forerunner token heads don’t show selectivity toward the semantic consistency between input text and labels simply merge the input text representations into the label tokens.
@@
Hidden States of Label Tokens are Joint Representations of Text Encodings and Label Semantics. Given the findings above, we probe the content of hidden states of label tokens $[y_{t}]$, i.e., how the text representation interacts with the original label semantics. We first train two centroid classifiers to predict the corresponding label $y_{t}$: (1) $\mathcal{C}_{f}$ trained on the hidden states of forerunner tokens $[s_{t}]$ and (2) $\mathcal{C}_{l}$ trained on the hidden states of label tokens $[y_{t}]$. To check whether the label tokens include the information of forerunner tokens, we use $\mathcal{C}_{f}$ to predict the label on the hidden state of label token $[y_{t}]$ in Fig. 5 (Right, solid). It shows that, during the copy processing, high classification accuracies can be achieved both on the correct label tokens and wrong label tokens, suggesting that the text features in the forerunner tokens can be partly and linearly detected in the label tokens. Moreover, results using $\mathcal{C}_{l}$ (dotted line) shows extreme results, suggesting the label information remains in the label token. So, we can conclude that: hidden states of label tokens are joint representations of label semantics and text representations.
@@
Label Denoising is Conducted on the Overlap of Label Semantics and Text Representations. Notice that in Fig. 5 (Right, solid), compared to the typical results predicted on the forerunner tokens, accuracies are improved on the correct label and suppressed on the wrong label, which suggests that information consistent with the label semantics is easier to be enhanced by the label tokens and vice versa, showing a feature selectivity on the consistency between the text representations and the label semantics. Given the observation that the information on label semantics and text features can be extracted separately and linearly, we can confirm that these two kinds of information are located in different sub-spaces of the hidden states, and linearly merged by the attention operation of forerunner heads. Moreover, given the fact that there is no selectivity is observed in the copy behavior of the forerunner token head (Fig. 5 (Middle)), it is intuitive that the feature selectivity shown in Fig. 5 (Right) comes from the arithmetical interaction of feature vectors on the overlap of sub-spaces between the label semantics and text features, making ICL stable against label noise (Min et al., 2022). Moreover, as mentioned by Wei et al. (2023), large models show poorer stability against label noise. We infer that the larger hidden dimensions in larger models lower the overlap between the sub-spaces of label semantics and text representations to reduce the interaction.
@@
Figure 5: Hidden states copy magnitude from forerunner tokens to label tokens against layers. Left: Kernel alignment between the forerunner token (the copy source) and the label token of the next layer (the copy target). Middle: Curves: The count of marked forerunner token heads with correct and wrong labels; Colored Areas: The maximum attention scores from forerunner token to query (copy magnitude) with correct and wrong labels (detailed attention head statistical data is in Appendix F.1). Right: Centroid classifier results predicted on the hidden states of correct and wrong label tokens, on SST-2 and MR. (Solid: Predicted by classifiers $\mathcal{C}_{f}$ trained on hidden states of forerunners. Dotted: Predicted by classifiers $\mathcal{C}_{l}$ trained on hidden states of label tokens.)

### v2

@@ Abstract
In-context Learning (ICL) is an emerging few-shot learning paradigm on Language Models (LMs) with inner mechanisms un-explored. There are already existing works describing the inner processing of ICL, while they struggle to capture all the inference phenomena in large language models. Therefore, this paper proposes a comprehensive circuit to model the inference dynamics and try to explain the observed phenomena of ICL. In detail, we divide ICL inference into 3 major operations: (1) Input Text Encode: LMs encode every input text (demonstrations and queries) into linear representation in the hidden states with sufficient information to solve ICL tasks. (2) Semantics Merge: LMs merge the encoded representations of demonstrations with their corresponding label tokens to produce joint representations of labels and demonstrations. (3) Feature Retrieval and Copy: LMs search the joint representations similar to the query representation on a task subspace, and copy the searched representations into the query. Then, language model heads capture these copied label representations to a certain extent and decode them into predicted labels. The proposed inference circuit successfully captured many phenomena observed during the ICL process, making it a comprehensive and practical explanation of the ICL inference process. Moreover, ablation analysis by disabling the proposed steps seriously damages the ICL performance, suggesting the proposed inference circuit is a dominating mechanism. Additionally, we confirm and list some bypass mechanisms that solve ICL tasks in parallel with the proposed circuit.
@@ 1 Introduction
In-Context Learning (ICL) (Radford et al., 2019; Dong et al., 2022) is an emerging few-shot learning paradigm: given the demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ consisting of [input text]-[label token] pairs and a query $x_{q}$, Language Models (LMs) take the sequence $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$ [^1] In this paper, we denote tokenization as []delimited-[][$\cdot$][ ⋅ ], and token concatenating as [][]delimited-[]delimited-[][$\cdot$][$\cdot$][ ⋅ ] [ ⋅ ]. (Fig. 1) as input and then predicts the label for $x_{q}$ by causal language modeling operation. Typically, the label tokens $y_{i}$ are preceded by and also predicted by forerunner tokens $s_{i}$ (e.g., the colon in “Label: ”). ICL has aroused widespread interest, but its underlying mechanism is still unclear.
@@
As shown in Fig. 1, we decompose ICL dynamics into 3 atomic operations on Transformer layers. Step 1: Input Text Encode: LMs encode each input text $x_{i}$ into linear representations in the hidden state of its corresponding forerunner token $s_{i}$. Step 2: Semantics Merge: For demonstrations, LMs merge the encoded representations of $s_{i}$ with the hidden state of its corresponding label tokens $y_{i}$. Step 3: Feature Retrieval and Copy: LMs retrieve merged label representations $y_{1:k}$ from Step 2 similar to the query representation $s_{q}$ in a task-relevant subspace and then merge them with the query representation. Finally, LM heads predict the label for $x_{q}$ using the label-attached query representation $s_{q}$. Steps 2 and 3 form a typical induction circuit, which is a key mechanism of ICL but only examined in synthetic scenarios (Elhage et al., 2021; Singh et al., 2024b; Reddy, 2024).
@@
We empirically find evidence for the existence of each proposed step in LLMs, and conduct more fine-grained measurements to gain insights into some phenomena observed in ICL scenarios, such as (1) positional bias: the prediction is more influenced by the latter demonstration (Zhao et al., 2021), (2) noise robustness: the prediction is not easy to be affected by demonstrations with noisy labels (Min et al., 2022), while larger models are less robust to noise (Wei et al., 2023), and (3) demonstration saturation: the accuracy improvements plateau when sufficient demonstrations are given (Agarwal et al., 2024; Bertsch et al., 2024), etc. (discussed in §5.3). Moreover, we find multiple bypass mechanisms for ICL with the help of residual connections, while the 3-phase dynamics remains dominant.
@@ 2.2 Experiment Settings
Models. We mainly conduct experiments on 4 modern LMs: Llama 3 (8B, 70B) (AI@Meta, 2024), and Falcon (7B, 40B) (Almazrouei et al., 2023). Unless specified, we report the results on Llama 3 70B, since its deep and narrow structure (80 layers, 64 heads) makes it easier to show hierarchical inference dynamics (discussed in §5.2). The results of other models can be found in Appendix H.2.
Datasets. We build ICL-formed test inputs from 6 sentence classification datasets, and unless specified, we report the average results on them: SST-2 (Socher et al., 2013), MR (Pang & Lee, 2005), Financial Phrasebank (Malo et al., 2014), SST-5 (Socher et al., 2013), TREC (Li & Roth, 2002; Hovy et al., 2001), and AGNews (Zhang et al., 2015).
Others. Unless specified, we use $k=4$ demonstrations in ICL inputs. For each dataset, we randomly sample $512$ test data and assign one order-fixed demonstration sequence for each test sample. About the prompt templates, etc., please refer to Appendix A.1.
3 Step 1, Input Text Encode: Semantics Encoding as Linear Representations in Hidden States
@@
3.1 LLMs Encode Input Text on Forerunner Tokens in Hidden States
We first study the existence of input text encoding in hidden states and then explain their linear separability and task relevance in §3.2. For each text-label pair $(x_{t},y_{t})$ (encoding target) sampled from the datasets, we prepend them with $k$ demonstrations, resulting in augmentated (by label $[y_{t}]$) ICL-styled inputs $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{t}][s_{t}][y_{t}]$. These inputs are then fed into an LM to extract the hidden states of a specific token in $[x_{t}][s_{t}][y_{t}]$ from each layer, serving as the ICL inner representations. To assess the quality of these representations as sentence representations, we use the sentence embedding of $x_{t}$ encoded by BGE M3 (Chen et al., 2024), a SotA encoder-only Transformer, as a reference representation and then calculate the mutual nearest-neighbor kernel alignment [^3] Intuitively, kernel alignment measures similarity between two representations toward the same datasets, and according to Huh et al. (2024), a higher cross-model kernel alignment usually means a better representation. (Huh et al., 2024) between these representations. See Appendix A.1 and A.2 for details.
@@ 3.1 LLMs Encode Input Text on Forerunner Tokens in Hidden States
Perplexed Texts are Encoded Worse. We investigate the correlation between kernel alignment and the perplexity of encoding targets with different $k$. Fig. 2 (Right, upper) shows a negative correlation for $k=0$, that is, LMs generate poorer encodings for more complex input text when no demonstrations are given, which can be identified as an In-weight Learning (IWL) property of the inner text encoding. While, when demonstrations are given in context (Fig. 2 (Right, lower)), the negative correlation disappears, which suggests that LMs effectively encode more complex samples with the help of demonstrations in context. More discussion about the relationship between classification performance and perplexity is in Appendix F.
@@
3.2 Input Text Encoding is Linear and Task-Relevant but Position-biased
Figure 2: Input text encoding magnitudes (metricized by kernel alignment against feature encoded by an encoder-structured model) of hidden states in various layers in ICL scenario (The controlled experiments are results between current 6 datasets and TEE (Mohammad et al., 2018)). Left: Encoding magnitudes on hidden states from various types of token. Middle [^5] Experiments of Fig. 2 (Middle) on Llama 3 70B do not involve results on AGNews.: Encoding magnitudes with different $k$ on the forerunner tokens. Right: Encoding magnitudes in layer 24 of Llama 3 70B against the causal language modeling loss of the input text with (upper) $k=0$ and (lower) $k=8$.
@@ 3.2 Input Text Encoding is Linear and Task-Relevant but Position-biased
4 Induction Circuits in Large Language Models
Figure 4: The similarities of ICL hidden states in different positions on layer 24 between Left: the same queries, Right: two different queries (on SST-2).
This section mainly shows how LMs utilize the encoded linear text representations in induction circuits with a typical 2-step form (Singh et al., 2024b): Forerunner Token Heads merge the demonstration text representations in the forerunner token into their corresponding label tokens with a selectivity regarding the compatibility of demonstrations and label semantics. Induction Heads copy the information in the label representations similar to the query representations back to the query. Such operations are conducted on task-specific subspaces, enabling LMs to solve multiple tasks by multiplexing hidden spaces.
@@ 4.1 Step 2, Forerunner Token Head: Copy from Text Feature to Label Token
Text Representations are Copied without Selectivity. For each attention head, we extract the attention score [^7] This paper use notation $\alpha_{k\rightarrow q}$ to denote attention score with attention query $q$ and attention key $k$. $\alpha_{s_{t}\rightarrow y_{t}}$ from a forerunner token $[s_{t}]$ (as attention key) to the corresponding label token $[y_{t}]$ (as attention query). We then mark the head with $\alpha_{s_{t}\rightarrow y_{t}}\geqslant 5/n_{t}$ ($n_{t}$: the length of tokens before $[y_{t}]$) as a Forerunner Token Head and count them in each layer. The results are shown in Fig. 5 (Middle, “Correct Label”), where the peak matches the copy period in Fig. 5 (Left). Moreover, to investigate the influence of the correctness of label tokens, we replace $[y_{t}]$ with a wrong label token [^8] For example, for a label space of “positive” and “negative”, if $[y_{t}]$ is “positive”, we replace it to “negative”., where the results in Fig. 5 (Middle, “Wrong Label”) are almost identical to the correct-label setting, suggesting that the forerunner token heads don’t show selectivity toward the semantic consistency between input text and labels simply merge the input text representations into the label tokens. Furthermore, we find (Appendix D) that the copy processing is inherent: during the copy processing, LMs establish strong attention connections to preceding tokens, regardless of whether these copied tokens are forerunner tokens or not, which indicates that such copy processing is a universal inference behavior developed during pre-training, rather than being triggered by the special tokens in the in-context learning (ICL) input, while still aiding ICL processing (to be discussed in §5.1).
@@
Hidden States of Label Tokens are Joint Representations of Text Encodings and Label Semantics. Given the findings above, we probe the content of hidden states of label tokens $[y_{t}]$, i.e., how the copied text representation interacts with the original label semantics. We first train two centroid classifiers to predict the corresponding label $y_{t}$: (1) $\mathcal{C}_{f}$ trained on the hidden states of forerunner tokens $[s_{t}]$ and (2) $\mathcal{C}_{l}$ trained on the hidden states of label tokens $[y_{t}]$. To check whether the label tokens include the information of forerunner tokens, we use $\mathcal{C}_{f}$ to predict the label on the hidden state of label token $[y_{t}]$ in Fig. 5 (Right, solid). It shows that, during the copy processing, high classification accuracies can be achieved both on the correct label tokens and wrong label tokens, suggesting that the text features in the forerunner tokens can be partly and linearly detected in the label tokens. Moreover, results using $\mathcal{C}_{l}$ (dotted line) shows extreme results, suggesting the label information remains in the label token. So, we can conclude that: hidden states of label tokens are joint representations of label semantics and text representations. Moreover, interestingly, the accuracies from $\mathcal{C}_{f}$ shown in Fig. 5 (Right) decline after layer 35, suggesting that the information sharing between the forerunner tokens and the label tokens ends in later layers, which aligns with the results in Fig. 5 (Middle).
@@
Label Denoising is Conducted on the Overlap of Label Semantics and Text Representations. Notice that in Fig. 5 (Right, solid), compared to the typical results predicted on the forerunner tokens, accuracies are improved on the correct label and suppressed on the wrong label, which suggests that information consistent with the label semantics is easier to be enhanced by the label tokens and vice versa, showing a feature selectivity on the consistency between the text representations and the label semantics. Given the observation that the information on label semantics and text features can be extracted separately and linearly, we can confirm that these two kinds of information are located in different sub-spaces of the hidden states, and linearly merged by the attention operation of forerunner heads. Moreover, given the fact that there is no selectivity is observed in the copy behavior of the forerunner token head (Fig. 5 (Middle)), it is intuitive that the feature selectivity shown in Fig. 5 (Right) comes from the arithmetical interaction of feature vectors on the overlap of sub-spaces between the label semantics and text features, making ICL stable against label noise (Min et al., 2022). Moreover, as mentioned by Wei et al. (2023), large models show poorer stability against label noise, and we infer that the larger hidden dimensions in larger models lower the overlap between the sub-spaces of label semantics and text representations to reduce the interaction.
@@
Figure 5: Hidden states copy magnitude from forerunner tokens to label tokens against layers. Left: Kernel alignment between the forerunner token (the copy source) and the label token of the next layer (the copy target). Middle: Curves: The count of marked forerunner token heads with correct and wrong labels; Colored Areas: The maximum attention scores from forerunner token to query (copy magnitude) with correct and wrong labels (detailed attention head statistical data is in Appendix H.1). Right: Centroid classifier results predicted on the hidden states of correct and wrong label tokens, on SST-2 and MR. (Solid: Predicted by classifiers $\mathcal{C}_{f}$ trained on hidden states of forerunners. Dotted: Predicted by classifiers $\mathcal{C}_{l}$ trained on hidden states of label tokens.)